OpenStudy (anonymous):

A test is worth 100 points. There are 29 questions. Each question is worth either 2 or 5 points. How many of each value are on the test?

OpenStudy (jackellyn):

Let x=3pt questions and 29-x=5 point questions 2x+5(29-x)=100 Distribute and simplify 2x + 145 - 5x =100 -3x + 145 = 100 -3x = -45 x=15 So there are 15, 2pt questions and 29-15=14, 5 pt questions check: 15(2)+14(5)=100 100=100 @tr0nkitty